OpenStudy (anonymous):

the girl to boy ratio is 4 to 7. if there is 100 students total how many boys are there

OpenStudy (anonymous):

i don't think its 70

OpenStudy (acxbox22):

@kah.x if there are 70 boys and 30 girls then girl/boy ratio is 30/70=3/7\[\frac{ 3 }{ 7 }\neq \frac{ 4 }{ 7 }\]

OpenStudy (acxbox22):

it is wrong....

OpenStudy (anonymous):

i typed it on the wrong tab lol sorry

OpenStudy (acxbox22):

this question wouldnt make sense because the number of boys is not a whole number

OpenStudy (anonymous):

how do i figure out the answer ?

OpenStudy (anonymous):

acxbox22 is correct. if the ratio of girls to boys is 4:7, then 7/11 of the classroom is boys. 7 divided by 11 is 0.63636363... or 63.63%. 63.63% of 100 students is 63.63. with that being said there cant be 63.63 boys, so the answer is either 63 or 64, id say 64 because you should round up.